You have 100 seats in a theater, it cost $5 for each men to get in, it costs $2 for each women to get in and $0.50 for each kid to get in, you have $100. Using that whole $100 to fill up all the seats in the theater (100), how many men, how many women and how many kids can you buy tickets for?
Answer by ewatrrr(24785)

Re TY
Now understand Your Question. Ty
It is of the variety ...
ax + by + cz = d
x + y + z = 100 (seats)
....
5x + 2y + .5z = $100
$5(k) + $2(100-k-m) + $0.5m = $100
solve by assigning appropriate values to k & m
Using
k = 5 , m = 70
$5(5) + $2(25) + $0.5(70) = $100
Ans: 5M, 25W and 70 children